OpenStudy (anonymous):

What is the standard form equation of the line passing through (5, –3) with a slope of –3?

OpenStudy (anonymous):

Equation will be: \[\huge \color{green}{y - y_1 = m(x - x_1)}\] m= -3 \(x_1, y_1\) = (5, -3) Plug in the values and then solve..

OpenStudy (anonymous):

After plugging in the values you will get: \[y +3 = -3(x - 5)\] \[y + 3 = -3x + 15\] \[y = -3x + 12\]

OpenStudy (anonymous):

thank you!